# Exploring fully-heavy tetraquarks through the CGAN framework: Mass and width

### v1

@@
M. Malekhosseini1 , S. Rostami1, A. R. Olamaei2,3, K. Azizi1,4
@@ 3 School of Physics, Institute for Research in Fundamental Sciences (IPM),
4 Department of Physics, Doğuş University, Dudullu-Ümraniye, 34775 Istanbul, Türkiye
@@ 3.1 $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ states
Table 1: Our CGAN predictions for the mass of the $cc\bar{c}\bar{c}$ state (in units of MeV), based on A1 and A2, compared with the experimental data [3].
@@
Table 2: Our CGAN predictions for the full width of the $cc\bar{c}\bar{c}$ state (in units of MeV), based on A1 and A2, compared with the experimental data [3].
@@ 3.5 The $cb\bar{c}\bar{b}$ state
Finally, we examine the $cb\bar{c}\bar{b}$ system. This system is neutral and has a certain C-parity. Table 11 presents mass predictions of the CGAN model (A1 and A2) for the $cb\bar{c}\bar{b}$ system across various quantum states ($J^{PC}$), compared to multiple theoretical models. It is obvious that the predicted results of the CGAN model based on A2 tend to be slightly higher compared to A1, indicating that the change in dataset features between these two approaches influences the predictions. However, the CGAN results, based on both approaches, are in good agreement with the theoretical model results, with only minor discrepancies. Table 12 illustrates the width predictions for the $cb\bar{c}\bar{b}$ state from our CGAN model (A1 and A2) in comparison with the QCD sum rule method. The width predictions from the CGAN model based on A2 are generally higher than those from A1. When compared to the available QCD sum rule results, the CGAN predictions show a trend of being smaller than the QCD sum rule values. For instance, the mass spectra of fully-heavy tetraquarks predicted by the CGAN model based on A2 for various quantum states are shown in Fig. 1, demonstrating the capabilities of this method in exploring heavy tetraquark systems.

### v2

@@
M. Malekhosseini1 , S. Rostami2, A. R. Olamaei2,3, K. Azizi1,4
@@ 2 School of Physics, Institute for Research in Fundamental Sciences (IPM),
4 Department of Physics, Dogus University, Dudullu-Ümraniye, 34775 Istanbul, Türkiye
@@ 3.1 $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ states
The QCD sum rule method Ref. [61] 6540 - - 6520 7000 6980 - - 6470 - 7000 6990 -
3
Ref. [39, 62] 5990 - 6090 - - - - 6050 - - 6110 -
Ref. [26] 6570 - - - - - - - - - - -
The nonrelativistic constituent quark models Ref. [34] 6384 6512 - 6483 - - - - 6452 - - - -
3
Ref. [63] 6377 6425 - 6432 - - - - 6425 - - - -
6371 6483 - 6479 - - - - 6450 - - - -
3
Ref. [15] 6487 6518 - 6524 - - - - 6500 - - - -
3
Ref. [64] 6500 6411 - 6475 - - - - 6453 - - - -
3
The Bethe- Salpeter equations Ref.[65] 6419 - 6516 - - - - 6456 - - - -
The relativistic quark model Ref. [66] 6190 - 6367 - - - - 6271 - - - -
3
Ref. [67] 6435 6542 - 6543 - - - - 6515 - - - -
Monte Carlo method Ref. [30] 6351 - 6471 - - - - 6441 - - - -
The diquark model Ref. [24] 5966 - 6223 - - - - 6051 - - - -
3
Ref. [68] 6322 - 6385 - - - - 6354 - - - -
An effective potential model Ref. [69] 6346 6476 - 6475 - - - - 6441 - - - -
Multiquark color flux-tube model Ref. [31] 6407 - 6486 - - - - 6463 - - - -
The chromo- magnetic model Ref. [25] 6192 - - - - - - - - - - -
3
Ref. [70] 6899 7016 - 6956 - - - - 6899 - - - -
6035 6253 - 6194 - - - - 6137 - - - -
3
Ref. [71] 6045 6271 - 6287 - - - - 6231 - - - -
3
Ref. [72] 6034 6254 - 6194 - - - - 6137 - - - -
Table 1: Our CGAN predictions for the mass of the $cc\bar{c}\bar{c}$ state (in units of MeV), based on A1 and A2, compared with the experimental data [3] and the theoretical estimates.
@@
The QCD sum rule method Ref. [26] 110 - - - - - - - - - - -
3
Ref. [27] 128 - - - - - - - - - - -
3
Ref. [73] 70.18 - - - - - - - - - - -
Table 2: Our CGAN predictions for the full width of the $cc\bar{c}\bar{c}$ state (in units of MeV), based on A1 and A2, compared with the experimental data [3] and the theoretical estimates.
@@ 3.5 The $cb\bar{c}\bar{b}$ state
Finally, we examine the $cb\bar{c}\bar{b}$ system. This system is neutral and has a certain C-parity. Table 11 presents mass predictions of the CGAN model (A1 and A2) for the $cb\bar{c}\bar{b}$ system across various quantum states ($J^{PC}$), compared to multiple theoretical models. It is obvious that the predicted results of the CGAN model based on A2 tend to be slightly higher compared to A1, indicating that the change in dataset features between these two approaches influences the predictions. However, the CGAN results, based on both approaches, are in good agreement with the theoretical model results, with only minor discrepancies. Table 12 illustrates the width predictions for the $cb\bar{c}\bar{b}$ state from our CGAN model (A1 and A2) in comparison with the QCD sum rule method. The width predictions from the CGAN model based on A2 are generally higher than those from A1. When compared to the available QCD sum rule results, the CGAN predictions show a trend of being smaller than the QCD sum rule values. For instance, the mass spectra of fully-heavy tetraquarks predicted by the CGAN model based on A2 for various quantum states are shown in Fig. 1, demonstrating the capabilities of this method in exploring heavy tetraquark systems. It is worth noting that, unlike traditional quark models where the mass hierarchy typically increases with total angular momentum $J$, our predictions focus on reproducing the experimentally observed mass hierarchy and determining which angular momentum value leads to a mass prediction that most closely matches the experimental result (specifically for $cc\bar{c}\bar{c}$) or presents a more reasonable result (for other systems), rather than examining a spectrum of states with identical quark content, where a systematic increase in mass with spin is typically observed. In other words, our approach focuses on identifying the most probable spin configuration corresponding to the known mass, even if the predicted hierarchy does not strictly follow conventional expectations.